40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 21 16:50:30 2012, Rana, Update, SUS, TT in vac DB25 pin swapping 
    Reply  Wed Jan 2 06:52:50 2013, Koji, Update, SUS, TT in vac DB25 pin swapping 
Message ID: 7871     Entry time: Wed Jan 2 06:52:50 2013     In reply to: 7869
Author: Koji 
Type: Update 
Category: SUS 
Subject: TT in vac DB25 pin swapping 

[Koji, Rana, Nic, Steve]

I recalled that we used an optical lever to check if the TT is moving or not.
We used a laser pointer on a tripod, which was prepared by Steve.

I should also note that we stepped back the eddy current dampers from the magnets
in order to enhance the motion of the suspension. They should be restored in the end.

The mini-D connectors on the OSEMs are loosened so that we can plug the cables in.
This requires a specific metric allen key
that is stored in a clean tool box with an aluminum foil.

ELOG V3.1.3-